OpenStudy (anonymous):

Solve the equation. Be sure to show your work. All final answers must be in simplest form. P+2/3=6

OpenStudy (mertsj):

\[P+\frac{2}{3}=6\]

OpenStudy (mertsj):

Subtract 2/3 from both sides. Can you do that?

OpenStudy (y2o2):

p = 6-(2/3) p = 16/3

OpenStudy (anonymous):

Thanks everyone

OpenStudy (mertsj):

yw